### v1

@@ 2.2 Universal Approximation Theorem
Hua et al. (2023) [14] treat a general, possibly nonlinear, operator
@@
In this work, we restrict ourselves to the transfer and Koopman operators associated with a measure‑preserving dynamical system. Under these conditions, both operators satisfy
$\mathcal{L}\colon L^{2}(X)\to L^{2}(X),$
so only a single subspace $V\subset L^{2}(X)$ is required to approximate both the observable $g$ and its image ^$\hat{\mathcal{L}}(g)$. Consequently, one should instead use a single encoder. We therefore state our approximation theorem in this alternate setting.
@@
Let $X\subset\mathbb{R}^{d}$ be compact and let $\mathcal{L}:L^{2}(X)\rightarrow L^{2}(X)$ be a continuous operator that is either the transfer or the Koopman operator associated with some measure-preserving dynamical system on $X$. Then for each compact set $\mathcal{K}\subset L^{2}(X)$ and any $10<\epsilon<$ 1 there exist $N>0$, a neural basis $\phi=(\phi_{1},\phi_{2},\dots,\phi_{N}):X\rightarrow\mathbb{R}^{N}$ and a linear map $\mathcal{G}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ such that the operator $A\coloneq\mathcal{D}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{G}\circ\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}$ satisfies
$\sup_{g\in\mathcal{K}}\left\|\mathcal{L}(g)-A(g)\right\|_{L^{2}(X)}\leq\epsilon.$
@@
Typical examples of compact sets $\mathcal{K}\in L^{2}(X)$, with $X\subset\mathbb{R}^{b}$ ($b\in\mathbb{N}$), include the closed unit ball of any finite-dimensional subspace $\operatorname{span}\{\chi_{1},\dots,\chi_{m}\}\subset L^{2}(X)$ and - when $X$ is a bounded Lipschitz domain (its boundary is locally a Lipschitz graph) - the closed unit ball of $H^{1}(X)$; the latter is compact in $L^{2}(X)$ by the Rellich–Kondrachov embedding
$H^{1}(X)\hookrightarrow L^{p^{\ast}}(X),\qquad p^{\ast}=\frac{2b}{\,b-2\,},$
valid for suitable $b$, and for either choice of $\mathcal{K}$ one can choose a radius $R>0$ so that the spatial and Fourier integrals in [23, Theorem 3] are arbitrarily small.
Proof 2.2 (Proof of Theorem 2.1).
@@
The proof proceeds almost exactly as in [14, Theorem 2.2], so we omit the repeated details and focus on the necessary modifications. First, we need only one neural basis ($(\{\phi_{1},\ldots,\phi_{N}\}\subset L^{2}(X))$ for both the input and output spaces. The universal approximation property [14, Theorem 2.1] ensures we can approximate each $g\in\mathcal{K}$ by a finite‐rank projection $g^{N}$, and the continuity of $\mathcal{L}$ on an augmented compact set still supports a uniform‐continuity argument just as in the original proof.
@@
The main difference appears in the error decomposition. In [14], one bounds three terms to handle potentially different neural bases for input and output spaces, but here we merge everything into two terms. If $A\coloneq\mathcal{D}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{G}\circ\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}$, then for $g\in\mathcal{K}$
$\|\mathcal{L}(g)-A(g)\|_{L^{2}(X)}$ $\leq\Bigl{\|}\mathcal{L}(g)-\mathcal{D}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{L}(g)\Bigr{\|}_{L^{2}(X)}$
$+\Bigl{\|}\mathcal{D}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{L}(g)-\mathcal{D}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{G}\circ\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}(g)\Bigr{\|}_{L^{2}(X)},$
where the first term is controlled by choosing $N$ sufficiently large so that $\mathcal{D}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}$ approximates the identity on the relevant compact set and the second term is made small by learning $\mathcal{G}$ to approximate $\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{L}(g)$ from the input $\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}(g)$. Everything else remains unchanged from [14, Theorem 2.2], and thus we conclude
$\sup_{g\in\mathcal{K}}\|\mathcal{L}(g)-A(g)\|_{L^{2}(X)}\leq\epsilon.$
@@ 3 Numerical Examples
In this section, we illustrate the results obtained with our architecture through two representative examples. We consider the Perron–Frobenius operator in both examples, but we can easily extend them to the Koopman operator. The first example considers the circle rotation, highlighting the recovery of spectral properties of the associated linear operator. The second example examines a nonlinearly perturbed version of Arnold’s cat map, demonstrating how the learned basis functions can adapt to the underlying system dynamics.
@@ 3.1 Data Generation
Both examples in this paper — the circle rotation and the perturbed cat map — are dynamical systems on the torus $\mathbb{T}^{d}$. We therefore require periodic observables, and we use random trigonometric polynomials to provide a rich class of training functions. To train the proposed architecture, we generate a dataset of $D$ random trigonometric polynomial functions defined on a $d$-dimensional domain. Each such function is constructed as a finite linear combination of sinusoidal basis functions (sine and cosine terms) in each dimension, with randomly chosen coefficients and frequencies. Formally, a general $d$-dimensional trigonometric polynomial of maximum order $K$ can be written as a linear combination of terms of the form $\prod_{j=1}^{d}\Psi_{j}(k_{j}x_{j})$, where each $\Psi_{j}(\cdot)$ is either $\sin(\cdot)$ or $\cos(\cdot)$, and each integer frequency $k_{j}$ ranges from $1$ to $K$ for $\sin(\cdot)$ and from 00 to $K$ for $\cos(\cdot)$. In other words, no oscillatory component in any dimension exceeds frequency $K$. The coefficients in the linear combination are real numbers drawn at random uniformly in $[-1,1)$. Note that the inclusion of $k_{j}=0$ allows constant (zero-frequency) components in each dimension as well. For example, in $d=2$ dimensions (with coordinates $x_{1}$ and $x_{2}$), such trigonometric polynomials take the form:
$f(x_{1},x_{2})$ $=$ $c_{1}\sin(k_{1}x_{1})+c_{2}\cos(k_{2}x_{2})+c_{3}\sin(k_{1}x_{1})\sin(k_{2}x_{2})+c_{4}\sin(k_{1}x_{1})\cos(k_{2}x_{2})$
$+$ $c_{5}\cos(k_{1}x_{1})\sin(k_{2}x_{2})+c_{6}\cos(k_{1}x_{1})\cos(k_{2}x_{2}),$
with $k_{1},k_{2}\in\{1,\ldots,K\}$ for the sine terms and $k_{1},k_{2}\in\{0,1,\ldots,K\}$ for the cosine terms. In higher dimensions $d$, a general term in the $d$-dimensional polynomial is a product of trigonometric functions each of the $d$ coordinates, with frequencies up to $K$ in each coordinate. Such functions contains a rich variety of mixed-frequency components across all dimensions.
@@ 3.2 Examples setup
We train on collections of functions as described in Section 3.1. The size of each dataset is summarized in Table 1, where we note the number of training functions, the maximum trigonometric polynomial order, how many functions go into the validation and test sets, and the resolution of the grid.
@@
Both the validation and test subsets are drawn from the same underlying class of functions as the training data generated in Section 3.1. The training set is used to fit the model parameters by minimising the defined loss. During training, we periodically evaluate the current model on the validation set. If the validation error starts to rise while the training error keeps falling, we halt training (early stopping) to prevent overfitting. We also explore alternative hyperparameter settings—e.g. learning rate, network depth, and regularisation strength—and retain the configuration that achieves the lowest validation error. Only after all training and hyperparameter selection are complete do we assess the model on the test set. This provides an unbiased estimate of the model’s performance on unseen functions, since the test data is not used to guide model selection.
For more complex dynamics, like the nonlinearly perturbed cat map, we employ larger architectures to capture the increased complexity. The specifics of each architecture, including the dimensions, activation functions, and the number of basis learned basis functions are shown in Table 2.
Table 2: Architecture specifics for the circle rotation and perturbed cat map. We list the encoder ($\mathcal{E}$) and approximation networks ($\mathcal{G}$), their hidden layer architectures (e.g. $5\times 512$), and activation functions. The dashes indicate that there are no hidden layers for the network $\mathcal{G}$ since it is learning a linear map in $\mathbb{R}^{N}$.
Example Encoder ($\mathcal{E}$) $\mathcal{G}$ N
@@
Circle rotation 5 $\times$ 512 ReLU - Identity 30
@@
3.3 Transfer Operator with Circle rotation
The circle rotation map $T:S_{1}\to S_{1}$ is a transformation in the circle given by $T(x)=x+\alpha\pmod{2\pi}$. For our numerical examples, we chose $\alpha=1$. To avoid discontinuities at the endpoints of the domain interval, we embed the unit-radius circle $S^{1}$ into $\mathbb{R}^{2}$, identifying each $x$ with $(\cos x,\sin x)\in S^{1}$.
@@ 3.3 Transfer Operator with Circle rotation
After parameter optimisation, we obtain both a set of basis functions spanning a subspace invariant to the transfer operator and an estimator to its action. Fig. 3 displays how the learned estimator performs on unseen data for various sparsity scenarios.
Figure 3: Performance of the learned operator approximation on unseen input data, for different levels of sparsity.
The corresponding quantitative results—mean relative $L^{2}$ errors averaged over the test set—are summarised in Table 3.
@@
No Sparsity ($\beta_{3}=0$) $6.41\times 10^{-4}$
Sparsity ($\beta_{3}=0.1$) $6.83\times 10^{-4}$
Sparsity ($\beta_{3}=0.2$) $6.38\times 10^{-4}$
Sparsity ($\beta_{3}=0.6$) $7.32\times 10^{-4}$
The approximately invariant subspace is spanned by the set of basis functions depicted in Fig. 4. Each row in the heatmaps corresponds to a different basis function, while the horizontal axis represents the grid over which these functions are evaluated. By introducing a sparsity penalty, these functions gain increasingly localised support, as is evident in the more concentrated regions appearing in the heatmaps with higher penalty values.
Figure 4: Basis functions obtained under different levels of sparsity penalty. Higher sparsity encourages localisation, which is reflected in more concentrated regions of support.
After discarding functions with negligible norms and normalising the remaining ones, we examine the Gram matrix, defined as
$G_{kj}=\langle\phi_{k},\phi_{j}\rangle,\qquad k,j=1,\dots,N,$
in Fig. 5. The near-zero off-diagonal entries reveal that the introduction of the sparsity penalty moves the set of basis functions closer to orthogonality.
Figure 5: Gram matrix of the filtered and normalised basis functions, highlighting that the introduction of sparsity promotes near-orthogonality (off-diagonal elements remain close to zero).
We now highlight the recovery of the operator’s spectral properties through the learned estimator. By projecting onto the subspace spanned by $\{\phi_{j}\}_{j=1}^{N}$, we obtain the Galerkin system
$L_{kj}\;=\;\langle\phi_{k},\;\mathcal{G}(\phi_{j})\rangle\quad\text{and}\quad M_{kj}\;=\;\langle\phi_{k},\;\phi_{j}\rangle,$
where $M$ is often called the mass matrix. The approximate eigenpairs of $\hat{\mathcal{L}}$ are found by solving the generalized eigenvalue problem
$L\,\mathbf{v}=\lambda\,M\,\mathbf{v}.$
Eigenvalues and eigenvectors of $(L,M)$ then serve as approximations to the operator’s eigenpairs.
@@
Using the ansatz $\psi_{k}(\theta)=e^{ik\theta}$, we get
$(\mathcal{L}\,\psi_{k})(\theta)=e^{ik(\theta-\alpha)}=e^{-ik\alpha}\,\psi_{k}(\theta)\quad\Longrightarrow\quad\lambda_{k}=e^{-ik\alpha},\quad\psi_{k}(\theta)=e^{ik\theta}$
Thus, the leading eigenvalues take the form
$\lambda_{\pm 1}=e^{\mp i\alpha}.$
These analytic expressions match well with our numerical experiments, which were performed with the model parameter $\beta_{3}=0.6$.
Figure 6: Left: Eigenvalues of the Perron–Frobenius operator for a circle rotation by $\alpha=1$ radians. Numerical eigenvalues (blue; leading pair corresponding to $k=\pm 1$ in red) coincide with the analytic spectrum (green). Right: Magnitude versus argument of the same eigenvalues. The dashed green vertical lines mark angles that are integer multiples of $\alpha$, underscoring that the additional numerical eigenvalues are dynamically meaningful, but duplicate information from the lowest-order eigenvalues.
We reconstruct each leading eigenfunction from the Galerkin eigenvector $\mathbf{v}=(v_{1},\dots,v_{N})$—whose entries are the coefficients in our learned basis $\{\phi_{j}\}_{j=1}^{N}$—as
$\psi(\theta)=\sum_{j=1}^{N}v_{j}\,\phi_{j}(\theta),$
and find that it coincides perfectly with the analytic eigenfunction.
Figure 7: Real parts of the leading eigenfunctions of the transfer operator. The analytic eigenfunction is plotted in dashed black, while its numerical counterpart in solid red. In both cases the numerical eigenfunctions coincide with the analytic curves.
Finally, we demonstrate the invariance of the learned subspace under the Perron–Frobenius operator. Recall $A\coloneq\mathcal{D}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{G}\circ\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}$ is the data-driven approximation of the Perron-Frobenius operator. On the grid $\{x_{i}\}_{i=1,\dots,n}$ define
$\Phi=[\phi_{j}(x_{i})]_{i=1,\dots,n;\,j=1,\dots,N}\in\mathbb{R}^{n\times N},\qquad\Psi=[A(\phi_{j})(x_{i})]_{i=1,\dots,n;\,j=1,\dots,N}\in\mathbb{R}^{n\times N},$
and set $r=\operatorname{rank}\Phi=\operatorname{rank}\Psi\leq N$. We wish to determine how close the $r$-dimensional column space of $\Phi$ is to the $r$-dimensional column space of $\Psi$, and we do this by computing principal angles between these two subspaces. Following [12, Algorithm 6.4.3] we form the QR decomposition

### v2

@@ 2.2 Universal approximation theorem
Hua et al. (2023) [16] consider a general, possibly nonlinear, operator $\mathcal{J}$ between two distinct Hilbert spaces.
@@
In this work, we restrict ourselves to the transfer and Koopman operators $\mathcal{L}$ associated with a dynamical system. Each of these operators map $L^{2}(X)$ into $L^{2}(X)$, so only a single subspace $V\subset L^{2}(X)$ is required to approximate both the observable $g$ and its image $\mathcal{L}(g)$. Consequently, in contrast to [16], one should instead use a single encoder. We state an approximation theorem in this alternate setting.
@@
Let $X\subset\mathbb{R}^{d}$ be compact and let $T:X\to X$ be the evolution dynamics on $X$. Let $\mathcal{L}:L^{2}(X)\rightarrow L^{2}(X)$ be continuous and denote either the Koopman or the transfer operator associated with $T$. Given $N\in\mathbb{N}$ and a family $\phi=(\phi_{1},\dots,\phi_{N})\subset L^{2}(X)$ define
$\mathcal{P}_{\phi}\colon L^{2}(X)\to\mathbb{R}^{N},$ $\mathcal{P}_{\phi}(g)=\bigl{(}\langle g,\phi_{1}\rangle,\dots,\langle g,\phi_{N}\rangle\bigr{)}^{\!\top},$
$\mathcal{R}_{\phi}\colon\mathbb{R}^{N}\to L^{2}(X),$ $\mathcal{R}_{\phi}(a_{1},\dots,a_{N})=\sum_{j=1}^{N}a_{j}\phi_{j}.$
@@
Then for each compact set $\mathcal{K}\subset L^{2}(X)$ and any $0<\epsilon<1$ there exist $N>0$, a neural basis $\phi=(\phi_{1},\phi_{2},\dots,\phi_{N}):X\rightarrow\mathbb{R}^{N}$ and a linear map $\mathcal{G}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ such that the operator ^$\hat{\mathcal{L}}\coloneq\mathcal{R}_{\phi}\circ\mathcal{G}\circ\mathcal{P}_{\phi}$ satisfies
$\sup_{g\in\mathcal{K}}\left\|\mathcal{L}(g)-\hat{\mathcal{L}}(g)\right\|_{L^{2}(X)}\leq\epsilon.$
Remark 2.2.
@@
Examples of compact sets $\mathcal{K}\in L^{2}(X)$ with $X\subset\mathbb{R}^{d}$, include (i) the closed unit ball of any finite-dimensional subspace in $L^{2}(X)$ and (ii) the closed unit ball of $H^{1}(X)$, for $d\geq 2$ and $X$ a bounded $C^{1}$ domain (its boundary is locally a $C^{1}$ graph). The latter is compact in $L^{2}(X)$ by the Rellich–Kondrachov Embedding Theorem (e.g. Theorem 9.16 [6]).
Two convenient sufficient conditions that guarantee the continuity of $\mathcal{L}$ are:
  1. 1.
    $T$ is measurable and $\mu-$invariant, so that $\mathcal{L}$ acts as an isometry on $L^{2}(X)$. The circle rotation in Section 3.3 is a representative example with $\mu=\mathrm{Lebesgue}$;
  2. 2.
    $T$ is non-singular with respect to $\mu$ and there exists a constant $c>0$ such that
    $0<c^{-1}\leq\dfrac{d(\mu\circ T^{-1})}{d\mu}\leq c<\infty$
    $\mu$ almost everywhere. The Anosov map in Section 3.4 is a representative example with $\mu=$ Lebesgue.
Proof 2.3 (Proof of Theorem 2.1).
@@
The proof proceeds almost exactly as in [16, Theorem 2.2], so we omit the repeated details and focus on the necessary modifications. We need only one neural basis $\{\phi_{1},\ldots,\phi_{N}\}\subset L^{2}(X)$ for both the input and output spaces. The universal approximation property [16, Theorem 2.1] ensures we can approximate each $g\in\mathcal{K}$ by an element $g^{N}\in\operatorname{span}\{\phi_{j}\}_{j=1}^{N}$. Let $\mathcal{K}_{\mathcal{L}}:=\mathcal{K}\cup\mathcal{L}(\mathcal{K})$ and choose $N$ and $\phi$ by [16, Theorem 2.1] so that, writing $Q:=\mathcal{R}_{\phi}\mathcal{P}_{\phi}$,
$\sup_{v\in\mathcal{K}_{\mathcal{L}}}\|v-Qv\|_{L^{2}(X)}\leq\eta\quad\text{for some }\eta>0.$
Next, since $Q$ is continuous and $\mathcal{K}$ is compact, $\mathcal{K}_{\mathcal{QL}}:=\mathcal{K}_{\mathcal{L}}\cup Q(\mathcal{K}_{\mathcal{L}})$ is compact, hence $\mathcal{L}$ is uniformly continuous on $\mathcal{K}_{\mathcal{QL}}$ with modulus of continuity $\omega_{\mathcal{L}}$.
@@
The main difference appears in the error decomposition. In [16], one bounds three terms to handle potentially different neural bases for input and output spaces, but here we merge everything into two terms. If ^$\hat{\mathcal{L}}\coloneq\mathcal{R}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{G}\circ\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}$, then for $g\in\mathcal{K}$
$\|\mathcal{L}(g)-\hat{\mathcal{L}}(g)\|_{L^{2}(X)}$ $\leq\Bigl{\|}\mathcal{L}(g)-\mathcal{R}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{L}(g)\Bigr{\|}_{L^{2}(X)}$
+$\quad+\Bigl{\|}\mathcal{R}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{L}(g)-\mathcal{R}_{\{\phi_{j}\}_{j=1}^{N}}\circ\mathcal{G}\circ\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}(g)\Bigr{\|}_{L^{2}(X)}.$
For the first term, since $\mathcal{L}(g)\in\mathcal{L}(\mathcal{K})\subset\mathcal{K}_{\mathcal{L}}$, we have $\|\mathcal{L}(g)-Q\mathcal{L}(g)\|\leq\eta$. Setting $\mathcal{G}:=\mathcal{P}_{\phi}\circ\mathcal{L}\circ\mathcal{R}_{\phi}$ in the second term we obtain
$\|Q\,\mathcal{L}(g)-Q\,\mathcal{L}(Qg)\|\leq\|Q\|\,\|\mathcal{L}(g)-\mathcal{L}(Qg)\|\leq\|Q\|\,\omega_{\mathcal{L}}(\|g-Qg\|)\leq\|Q\|\,\omega_{\mathcal{L}}(\eta).$
Everything else remains unchanged from [16, Theorem 2.2], and thus we conclude
$\sup_{g\in\mathcal{K}}\|\mathcal{L}(g)-\hat{\mathcal{L}}(g)\|_{L^{2}(X)}\leq\eta+\|Q\|\,\omega_{\mathcal{L}}(\eta).$
Choosing $N$ and $\phi$ so that the right-hand side is smaller than $\epsilon$ yields the claim.
@@ 3 Numerical examples
In this section, we illustrate the results obtained with our architecture through two representative examples. We consider the Perron–Frobenius operator in both examples, but the procedure is identical for training data generated by the Koopman operator. The first example considers a circle rotation, highlighting the recovery of spectral properties of the associated linear operator. The second example examines a nonlinearly perturbed version of Arnold’s cat map on the two-torus. This nonlinear map is highly anisotropic in state space and the eigenfunctions of the transfer operator are known to be distributions rather than functions [15, 9]. Through this very challenging example, we demonstrate how the learned basis functions can nevertheless successfully adapt to the system dynamics to construct an almost perfectly invariant subspace $V$.
@@ 3.1 Data generation
Both examples in this paper — the circle rotation and the perturbed cat map — are dynamical systems on periodic domains. We therefore require periodic observables, and we use random trigonometric polynomials to provide a rich class of training functions. To train the proposed architecture, we consider the intrinsic dimension $d^{\prime}=\dim\;X$ and generate a dataset of $D$ random trigonometric polynomial functions defined on a $d^{\prime}$-dimensional domain. Each such function is constructed as a finite linear combination of sinusoidal basis functions (sine and cosine terms) in each dimension, with randomly chosen coefficients. Formally, a general $d^{\prime}$-dimensional trigonometric polynomial of maximum order $K$ can be written as a linear combination of terms of the form $\prod_{j=1}^{d^{\prime}}\Psi_{j}(k_{j}x_{j})$, where each $\Psi_{j}(\cdot)$ is either $\sin(\cdot)$ or $\cos(\cdot)$, and each integer frequency $k_{j}$ ranges from 111 to $K$ for $\sin(\cdot)$ and from 0 to $K$ for $\cos(\cdot)$. In other words, no oscillatory component in any dimension exceeds frequency $K$. The coefficients in the linear combination are real numbers drawn at random uniformly in [1,1][-1,1][ - 1 , 1 ]. Note that the inclusion of $k_{j}=0$ allows constant (zero-frequency) components in each dimension as well. For example, in $d^{\prime}=2$ dimensions (with coordinates $x_{1}$ and $x_{2}$), such trigonometric polynomials take the form:
$f(x_{1},x_{2})=$ $c_{1}$ $\sin(k_{1}x_{1})\sin(k_{2}x_{2})+c_{2}\sin(k_{1}x_{1})\cos(k_{2}x_{2})$
$+$ $c_{3}\cos(k_{1}x_{1})\sin(k_{2}x_{2})+c_{4}\cos(k_{1}x_{1})\cos(k_{2}x_{2}),$
with $k_{1},k_{2}\in\{1,\ldots,K\}$ for the sine terms and $k_{1},k_{2}\in\{0,1,\ldots,K\}$ for the cosine terms. In the general $d^{\prime}$-dimensional case, each term is a product of trigonometric functions, one for each of the $d^{\prime}$ coordinates, with frequencies up to $K$ in each coordinate. Such functions contain a rich variety of mixed-frequency components across all dimensions.
@@ 3.2 Examples setup
We train on collections of functions as described in Section 3.1. The validation and test sets are drawn from the same underlying class of functions as our training data. The parameters of each dataset are summarized in Table 1.
@@
The training set is used to fit the model parameters by minimising the defined loss. During training, we periodically evaluate the current model on the validation set. We also explore alternative hyperparameter settings — e.g. learning rate, network depth, and regularisation strength—and retain the configuration that achieves the lowest validation error. Only after all training and hyperparameter selection are complete do we assess the model on the test set. This provides an unbiased estimate of the model’s performance on unseen functions, since the test data is not used to guide model selection.
For more complex dynamics, like the nonlinearly perturbed cat map, we employ larger architectures to capture the increased complexity. The specifics of each architecture, including the dimensions, activation functions, and the number of learned basis functions are shown in Table 2.
Table 2: Architecture specifics for the circle rotation and perturbed cat map. We list the encoder ($\mathcal{E}$) and approximation networks ($\mathcal{G}$), their hidden layer architectures (e.g. 5×5125$\times$ 5125 × 512 denotes five hidden layers with 512 units each), and activation functions. The dashes indicate that there are no hidden layers for the network $\mathcal{G}$ since it is learning a linear map in $\mathbb{R}^{N}$.
Example Encoder ($\mathcal{E}$) Approximation ($\mathcal{G}$) N
@@
Circle rotation 5 $\times$ 512 ReLU - Identity 19
@@
3.3 Transfer operator for a circle rotation
The circle rotation map $T:S^{1}\to S^{1}$ is a transformation in the circle given by $T(\theta)=\theta+\alpha\pmod{2\pi}$. For our numerical examples, we choose $\alpha=-1$. To avoid discontinuities at the endpoints of the domain interval, we embed the unit-radius circle $S^{1}$ into $\mathbb{R}^{2}$, identifying each $\theta$ with $(\cos\theta,\sin\theta)\in\mathbb{R}^{2}$. Thus the ambient dimension is $d=2$ while the intrinsic dimension remains $d^{\prime}=1$.
@@ 3.3 Transfer operator for a circle rotation
After the learning routine, we obtain both a set of basis functions spanning a subspace invariant under the transfer operator and an estimator for its action. Fig. 3 displays how the learned estimator performs on unseen data for various sparsity scenarios.
Figure 3: Performance of the learned operator approximation on unseen input data, for different levels of sparsity. No appreciable difference in performance is seen across the various sparsity penalties.
The corresponding quantitative results — mean relative $L^{2}$ errors averaged over the test set — are summarised in Table 3. Table 3 is consistent with Fig. 3; there is little change in relative error observed with increasing sparsity.
@@
No Sparsity ($\beta_{2}=0$) $3.974\times 10^{-3}$
Sparsity ($\beta_{2}=0.1$) $3.641\times 10^{-3}$
Sparsity ($\beta_{2}=0.2$) $3.958\times 10^{-3}$
Sparsity ($\beta_{2}=0.6$) $6.234\times 10^{-3}$
The approximately invariant subspace is spanned by the set of basis functions depicted in Fig. 4. By introducing a sparsity penalty, these functions gain increasingly localised support, as shown in Fig. 4.
Figure 4: Basis functions obtained under different sparsity penalties: each panel displays a distinct basis for a distinct sparsity penalty. In a single panel, the state space coordinates (the angles $\theta\in S^{1}$) are along the $x$-axis, and each row shows the values of a different basis function. Higher sparsity encourages localisation, which is reflected in more concentrated regions of support.
After $L^{2}$-normalising the learned basis, we examine the Gram matrix, defined as
(7) $M_{kj}=\langle\phi_{k},\phi_{j}\rangle,\qquad k,j=1,\dots,N,$
in Fig. 5. The introduction of a sparsity penalty slightly improves orthogonality.
Figure 5: Gram matrix of the filtered and normalised basis functions
We now recover the spectral properties of $\mathcal{L}$ from the learned estimator $\hat{\mathcal{L}}$. Consider the collection of learned basis functions $\{\phi_{j}\}_{j=1}^{N}\subset L^{2}(X)$. Suppose an observable $g\in L^{2}(X)$ can be represented as
$g=\sum_{j=1}^{N}\beta_{j}\phi_{j},$
where $\beta=(\beta_{1},\dots,\beta_{N})^{T}\in\mathbb{R}^{N}$ is the coefficient vector.
Applying the projection operator $\mathcal{P}_{\phi}$ to $g$ yields
$\mathcal{P}_{\phi}(g)=\begin{pmatrix}\langle g,\phi_{1}\rangle\\ \vdots\\ \langle g,\phi_{N}\rangle\end{pmatrix}=\begin{pmatrix}\langle\sum_{j=1}^{N}\beta_{j}\phi_{j},\phi_{1}\rangle\\ \vdots\\ \langle\sum_{j=1}^{N}\beta_{j}\phi_{j},\phi_{N}\rangle\end{pmatrix}=M\beta,$
where $M$ is the Gram matrix as defined in (7).
The learned operator is composed as ^$\hat{\mathcal{L}}=\mathcal{R}_{\phi}\circ\mathcal{G}\circ\mathcal{P}_{\phi}$ where $\mathcal{G}$ is an $N\times N$ matrix representing the dynamics in coefficient space. When $\hat{\mathcal{L}}$ is applied to $g$ we obtain the following linear transformation:
^$\hat{\mathcal{L}}g=\mathcal{R}_{\phi}(\mathcal{G}(\mathcal{P}_{\phi}(g)))=\mathcal{R}_{\phi}(\mathcal{G}(M\beta))=\sum_{j=1}^{N}(\mathcal{G}M\beta)_{j}\phi_{j}.$
A function $g$ with coefficient vector $\beta$ is an eigenfunction of $\hat{\mathcal{L}}$ with eigenvalue $\lambda$ if and only if ^$\hat{\mathcal{L}}g=\lambda g$. This means
$\sum_{j=1}^{N}(\mathcal{G}M\beta)_{j}\phi_{j}=\lambda\sum_{j=1}^{N}\beta_{j}\phi_{j}.$
Since the basis functions are approximately linearly independent, we must have
(8) $\mathcal{G}M\beta=\lambda\beta.$
Equation (8) is a generalised eigenvalue problem of size $N\times N$. Solving it yields eigenvalue-eigenvector pairs $(\lambda_{k},\beta^{(k)})$ for $k=1,\dots,N$. The corresponding eigenfunctions are
$g_{k}=\sum_{j=1}^{N}\beta_{j}^{(k)}\phi_{j},\quad k=1,\dots,N.$
@@
Using the eigenfunction ansatz $\psi_{k}(\theta)=e^{ik\theta}$ for $k\in\mathbb{Z}$, we have
$(\mathcal{L}\,\psi_{k})(\theta)=e^{ik(\theta-\alpha)}=e^{-ik\alpha}\,\psi_{k}(\theta)\quad\Longrightarrow\quad\lambda_{k}=e^{-ik\alpha},\quad\psi_{k}(\theta)=e^{ik\theta},$
which shows that we have eigenpairs $(\lambda_{k},\psi_{k}),k\in\mathbb{Z}$. Thus, the leading eigenvalues take the form
$\lambda_{\pm 1}=e^{\pm i\alpha}.$
These analytic expressions match well with our numerical experiments, which were performed with the model parameter $\beta_{2}=0.6$; see Fig. 6. Furthermore, the leading learned eigenfunctions coincide perfectly with the leading analytic eigenfunctions; see Fig. 7.
Figure 6: Left: Eigenvalues of the Perron–Frobenius operator for a circle rotation by $\alpha=-1$ radians. Numerical eigenvalues (blue circles; leading pair corresponding to $k=\pm 1$ shown as red circles) coincide with the analytic spectrum (green crosses) Right: Magnitude versus argument of the same eigenvalues. The dashed green vertical lines mark angles that are positive integer multiples of $\alpha$, underscoring that the additional numerical eigenvalues are dynamically meaningful, but duplicate information from the lowest-order eigenvalues.
Figure 7: Real parts of the leading eigenfunctions of the transfer operator. The analytic eigenfunction is plotted in dashed black, while its numerical counterpart is in solid red. In both cases the numerical eigenfunctions coincide with the analytic curves.
Finally, we demonstrate the invariance of the learned subspace under the Perron–Frobenius operator. On the grid $\{x_{i}\}_{i=1,\dots,n}$ define
$\Phi=[\phi_{j}(x_{i})]_{i=1,\dots,n;\,j=1,\dots,N}\in\mathbb{R}^{n\times N},\qquad\Psi=[\mathcal{L}(\phi_{j})(x_{i})]_{i=1,\dots,n;\,j=1,\dots,N}\in\mathbb{R}^{n\times N},$
and set $r=\operatorname{rank}\Phi=\operatorname{rank}\Psi\leq N$. We wish to determine how close the $r$-dimensional column space of $\Phi$ is to the $r$-dimensional column space of $\Psi$, and we do this by computing principal angles between these two subspaces. Following [14, Algorithm 6.4.3] we form the QR decomposition